A parallel circuit is set up With three resistors: 5 ohms 10 ohms and 15 ohms. If the power supply to the circuit is 30 V what is the current across each resistor and what is the total current through the circuit?


1
Expert's answer
2022-06-06T08:53:03-0400
"I_1=\frac{V}{R_1}=\frac{30\:\rm V}{5\:\Omega}=6\:\rm A"

"I_2=\frac{V}{R_2}=\frac{30\:\rm V}{10\:\Omega}=3\:\rm A"

"I_3=\frac{V}{R_3}=\frac{30\:\rm V}{15\:\Omega}=2\:\rm A"

"I_{\rm net}=I_1+I_2+I_3\\
=\rm 6\: A+3\: A+2\: A=11\: A"